OpenStudy (anonymous):

y=x^(1/3) surface area of revolution from y=1 to 2. edit: revolved around y axis

OpenStudy (anonymous):

\[2\pi \int\limits_{a}^{b}\sqrt{1+[f'x]^2}xdx\]

OpenStudy (anonymous):

f'x=.33x^-.666 [f'x]^2=.1111/x^1.3333 which is equal to 1/(9x^1.3333)

OpenStudy (sirm3d):

revolved about the x-axis or about the y-axis?

OpenStudy (anonymous):

y axis

OpenStudy (anonymous):

the equation i have should be right

OpenStudy (sirm3d):

\[\Large 2\pi\int_{y_1}^{y_2}x\sqrt{1+(x')^2}\;\mathrm dy\]

OpenStudy (sirm3d):

\[\Large y=x^{1/3}\Rightarrow y^3=x\]

OpenStudy (sirm3d):

it should be not difficult to integrate

OpenStudy (anonymous):

ok. for the x in the equation, do i just pu x or y^3

OpenStudy (sirm3d):

replace x by y^3

OpenStudy (anonymous):

o yay. i got it, thanks for the help.

OpenStudy (sirm3d):

yw